Part 1.

Supply Demand 10 20 30 40 50 60 70 80 Quantity

  1. What was the equilibrium price in this market before the tax?
  2. What is the amount of the tax?
  3. How much of the tax will the buyers pay?
  4. How much of the tax will the sellers pay?
  5. How much will the buyer pay for the product after the tax is imposed?
  6. How much will the seller receive after the tax is imposed?
  7. As a result of the tax, what has happened to the level of output?
  8. Calculate the economic welfare after government imposes a tax of $5 per unit on buyers.
    1. Total Surplus
    2. Government Revenue
    3. DWL
    4. Producer Surplus

Part 2.

Perfect Competition (Please Graph)

  1. Please explain and illustrate graphically how the diaper service market has been affected by the decrease in the North American birth rate and the development of disposable diaper.
  2. Explain the long-run and the short-run effects of the event, starting from the long run equilibrium.
  3. What happens to the price of diaper and the quantity of diaper in the market and a representative individual firm? (Show two diagrams for both market firms and an individual firm)
Answer #1

A.

As it can be seen in the figure that demand and supply is equal at price $8 and quantity 40 units.

Hence the equilibrium price in this market before the tax was $8.

B.

The amount of the tax=10-5

=$5

C.

The tax paid by buyers are =10-8

=$2

D

The tax paid by the sellers is=8-5

=$3